OpenStudy (anonymous):

What is the value of a in the equation 12a - 123 = 117? -1/2 20 2 228

OpenStudy (solomonzelman):

Add 123 to both sides Divide both sides by 12.

OpenStudy (nurali):

12a - 123 = 117 Add 123 and divide by 12.

OpenStudy (anonymous):

@Nurali So I would add 123 to 12 and 123?

OpenStudy (nurali):

Add 123 both side what will answer

OpenStudy (solomonzelman):

yes, add 123 to each of the sides.

OpenStudy (solomonzelman):

\[12a - 123\color{red}{ +123 } = 117\color{red}{ +123 }\]see how I am adding?

OpenStudy (nurali):

12a - 123 = 117 Add both side by 123 12a-123+123=117+123 12a=240 Divide by 12 a=240/12

OpenStudy (solomonzelman):

\[12a-123\color{red}{ +123 }=117\color{red}{ +123 }\]\[12a=240\]\[12a\color{red}{ \div 12 }=240\color{red}{ \div 12 }\]\[a~=~?\]

OpenStudy (anonymous):

a is 20. @SolomonZelman @Nurali

OpenStudy (nurali):

Correct

OpenStudy (solomonzelman):

YES!

OpenStudy (anonymous):

Thanks guys :) @Nurali @SolomonZelman

OpenStudy (solomonzelman):

YW!

OpenStudy (nurali):

My Pleasure.